5.Find the center and radius of the circle x^2+y^2+8x+6y=0
a.(âˆ′4,âˆ′3),r=5
b.(4,âˆ′3),r=5
c.(âˆ′4,3),r=5
d.None of the option

6.Find the equation of the line through the point (âˆ′1,2)which is parallel to y = 3x + 1
a.y=3x−5
b.y=−3x+5
c.y=−3x−5
d.y = 3x + 5
1
Expert's answer
2019-04-10T09:58:31-0400

5. x^2+y^2+8x+6y=0 -> (x+4)^2+(y+3)^2=25. Var. a.

6 If the line is parallel to y=2x+1, then y-2=2(x+1) -> y=2x+4.

If the line is parallel to y=3x+1, then y-2=3(x+1) -> y=3x+5. Var. d.